OpenStudy (anonymous):

how would you distribute x(x+3)= 14(x-4)?

OpenStudy (anonymous):

x*x=x^2 3*x=3x LHS of equation: x^2+3x 14*x=14x 14*(-4)=-56 RHS of equation: 14x-56 x^2+3x=14x-56

OpenStudy (anonymous):

how would you continue the problem?